Electricity Generated by Windmills:
One growing community in California currently uses 2 megawatts (MW) of electricity per year. All this electricity is generated from coal. The community plans to add x MW of windmill generation capacity to its system to meet increasing demand. How many megawatts of windmill generation capacity must be built in order that 20% of the community's total electricity usage comes from windmills?

I know the answer is supposed to be .5mw but I am stumped as to what formula to use to get that answer.
If I turn it into a sentence, I'd say "What is 20% of 2 MW?" That will give me a 0.4 answer though, which doesn't fit the correct answer. :(

Answer by jim_thompson5910(35256)

Present megawatt usage: 2 MW

Future megawatt usage: 2+x (we're adding on x megawatts)

-------------------------------------------------------

So we go from 2 to 2+x. The x megawatts added are coming from windmills and windmills alone (given in the instructions).


So that means we have...

(MW of windmills only)/(total megawatt usage) = percentage of windmill MW output

( x )/( 2 + x ) = 20%

( x )/( 2 + x ) = 20/100

( x )/( 2 + x ) = 0.20

x = 0.20( 2 + x )

x = 0.20(2) + 0.20(x)

x = 0.40 + 0.20x

x + 0.20x = 0.40

0.80x = 0.40

x = 0.40/0.80

x = 0.50

So you need to add 0.50 MW from windmills alone to get to 2+0.5=2.5 MW total

Notice how: 0.5/2.5 = 0.2 = 20%. So 20% of the energy is produced by windmills.